Liu Sales has two store locations. Sanford has fixed costs of $26,000 per month and a contribution margin ratio of 35%. Orlando has fixed costs of $440,000 per month and a contribution margin ratio of 65%. At what sales volume would the two stores have equal profits or losses?

a. cannot determine with the information given

b. $700,000

c. $600,000

d. $1,480,000
Answer #1

let the sale volume be x:

at this sale volume both the profits are equal:

profit = (sales * contribution margin ratio) - fixed cost

profit of sanford store = (x*0.35) - 26,000

profit of orlando store = (x*0.65)-440,000

at x both the profits are same.

(x*0.35)-26,000 = (x*0.65)-440,000

=>0.30x = 414,000

=>x =$1,380,000.

so at the sales volume of $1,380,000 both the stores will have equal profits.